# Analytic action principle in ${\cal N}=2$ AdS4 harmonic superspace

### v1

@@ Abstract
We develop the analytic action principle in the ${\cal N}=2$ AdS4 harmonic superspace and apply it for studying the component structure of the free $q$-hypermultiplet model.
1 Introduction and conclusions
Superspace formulations of supersymmetric field theories provide explicit off-shell realizations of the Poincaré superalgebra, whose closure does not require the equations of motion. Such formulations are particularly advantageous for models with extended supersymmetry, since hypermultiplet matter theories require an infinite number of auxiliary fields off shell, which naturally arise within the harmonic [1, 2] and projective [3, 4] superspace approaches. Being complementary [5], these two formalisms are especially useful for constructing classical actions of higher-spin field theories with extended supersymmetry [6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. Generalizing ${\cal N}=2$ supersymmetric field theory to the 4D anti de-Sitter background (AdS4), however, represents a nontrivial problem, as the corresponding geometry in the harmonic superspace has not yet been fully developed. Although general supergravity theories were formulated in 4D $\mathcal{N}=2$ harmonic superspace (see Refs. [1, 16, 17, 18, 19, 20, 21]), the specific features of ${\cal N}=2$ AdS4 harmonic superspace remain insufficiently explored, and the construction of classical actions for gauge and matter theories in these superspaces has not yet been properly addressed. In contrast, the analytic action principle in the AdS5 harmonic superspace was developed in the works [22, 23].
@@ 1 Introduction and conclusions
Despite the progress achieved in formulating field theories in ${\cal N}=2$ AdS4 harmonic superspace in the present paper and in Ref. [24], the problem remains far from complete. The approach of Ref. [24] is based on manifestly analytic superfields and employs the so-called analytic basis in superspace. In the present work, we use the central basis, in which analyticity is not manifest but is instead imposed through superfield constraints. This latter formulation is, however, fully covariant and explicitly geometric, owing to the developments in Refs. [25, 26, 27, 28, 29, 30, 31] where AdS4|8 superspace was obtained as a particular solution of $\mathcal{N}=2$ superfield supergravity. In general $\mathcal{N}=2$ supergravity, these two coordinate bases are related by the so-called bridge superfield, originally constructed in Ref. [20]. The explicit construction of this bridge superfield for the ${\cal N}=2$ AdS4 harmonic superspace background remains an important open problem, which will be addressed elsewhere.
The rest of the paper is organized as follows. In the next section, we present some background information about the geometry of the AdS4|8 superspace and its isometry generated by the Killing vectors. In Sec. 3 we extend this superspace with harmonic variables and introduce analytic superspace projectors. In Sec. 4 we show that these projectors may be used for construction a superfield action for an analytic Lagrangian in the harmonic superspace. In this section, we develop also the analytic action principle in a fully covariant and coordinate independent form. Then we apply this action principle for calculating the component structure of the classical action in the free $q$-hypermultiplet model. In our work, we follow superspace conventions adopted in Refs. [25, 26].
2 Isometries of AdS4|8 superspace
@@
$M_{ab},{\cal D}_{c}]=2\eta_{c[a}{\cal D}_{b]}\,,\quad[M_{\alpha\beta},{\cal D}^{i}_{\gamma}]=\varepsilon_{\gamma(\alpha}{\cal D}^{i}_{\beta)}\,,\quad[\bar{M}_{\dot{\alpha}\dot{\beta}},\bar{\cal D}^{i}_{\dot{\gamma}}]=\varepsilon_{\dot{\gamma}(\dot{\alpha}}\bar{\cal D}^{i}_{\dot{\beta})}\,.$ (2.3)
@@
$M_{ab}=(\sigma_{ab})^{\alpha\beta}M_{\alpha\beta}-(\tilde{\sigma}_{ab})^{\dot{\alpha}\dot{\beta}}\bar{M}_{\dot{\alpha}\dot{\beta}}\,,\quad M_{\alpha\beta}=\frac{1}{2}(\sigma^{ab})_{\alpha\beta}M_{ab}\,,\quad\bar{M}_{\dot{\alpha}\dot{\beta}}=-\frac{1}{2}(\tilde{\sigma}^{ab})_{\dot{\alpha}\dot{\beta}}M_{ab}\,.$ (2.4)
@@
$J_{kl},{\cal D}^{i}_{\alpha}]=-\delta^{i}_{(k}{\cal D}_{\alpha l)}\,,\qquad[J_{kl},\bar{\cal D}_{i}^{\dot{\alpha}}]=-\varepsilon_{i(k}\bar{\cal D}^{\dot{\alpha}}_{l)}\,.$ (2.6)
@@
$\{{\cal D}^{i}_{\alpha},{\cal D}^{j}_{\beta}\}=$ $4S^{ij}M_{\alpha\beta}+2\varepsilon_{\alpha\beta}\varepsilon^{ij}S^{kl}J_{kl}\,$, (2.7a)
$\{\bar{\cal D}_{\dot{\alpha}i},\bar{\cal D}_{\dot{\beta}j}\}=$ $-4S_{ij}\bar{M}_{\dot{\alpha}\dot{\beta}}+2\varepsilon_{\dot{\alpha}\dot{\beta}}\varepsilon_{ij}S^{kl}J_{kl}\,$, (2.7b)
$\{{\cal D}^{i}_{\alpha},\bar{\cal D}_{j}^{\dot{\beta}}\}=$ $-2i\delta^{i}_{j}(\sigma^{c})_{\alpha}{}^{\dot{\beta}}{\cal D}_{c}\,$, (2.7c)
$[{\cal D}_{a},{\cal D}^{j}_{\beta}]=$ $\frac{i}{2}(\sigma_{a})_{\beta\dot{\gamma}}S^{jk}\bar{\cal D}^{\dot{\gamma}}_{k}\,,\qquad[{\cal D}_{a},\bar{\cal D}_{j}^{\dot{\beta}}]=\frac{i}{2}(\sigma_{a})^{\alpha\dot{\beta}}S_{jk}{\cal D}^{k}_{\alpha}\,$, (2.7d)
$[{\cal D}_{a},{\cal D}_{b}]=$ $-S^{2}M_{ab}\,$, (2.7e)
@@
${\cal D}_{A}S^{ij}=0\,,\qquad\overline{S^{ij}}=S_{ij}\,.$ (2.8)
@@
$\xi=\xi^{A}{\cal D}_{A}=\xi^{a}{\cal D}_{a}+\xi^{\alpha}_{i}{\cal D}^{i}_{\alpha}+\bar{\xi}^{i}_{\dot{\alpha}}\bar{\cal D}_{i}^{\dot{\alpha}}$ (2.9)
@@
$\left[\xi+\frac{1}{2}\lambda^{ab}M_{ab}+\lambda^{ij}J_{ij},{\cal D}_{A}\right]=0\,$, (2.10)
for some real antisymmetric tensor $\lambda^{cd}(z)=-\lambda^{dc}(z)$ and real symmetric tensor $\lambda^{ij}=\overline{\lambda_{ij}}$. In Ref. [25] it was shown that the parameters $\lambda^{ij}$ may be expressed via a single superfield $\rho$, $\lambda^{ij}=\rho S^{ij}$. This superfield, as well as other components of the Killing vector $\xi^{A}$ obey a number of constraints which follow from Eq. (2.10):
${\cal D}^{i}_{\alpha}\xi^{\beta}_{j}-\rho S^{i}{}_{j}\delta_{\alpha}^{\beta}-\frac{1}{2}\lambda_{\alpha}{}^{\beta}\delta^{i}_{j}=$ $0\,$, (2.11a)
$\bar{\cal D}_{i}^{\dot{\alpha}}\xi^{\beta}_{j}-\frac{i}{2}S_{ij}\xi^{\beta\dot{\alpha}}=$ $0\,$, (2.11b)
$\bar{\cal D}_{i}^{\dot{\alpha}}\xi^{b}+2i(\sigma^{b})_{\beta}{}^{\dot{\alpha}}\xi_{i}^{\beta}=$ $0\,$, (2.11c)
${\cal D}^{i}_{\alpha}\lambda^{cd}-4S^{ij}\xi_{j}^{\beta}(\sigma^{cd})_{\alpha\beta}=$ $0\,$, (2.11d)
${\cal D}^{i}_{\alpha}\rho-2\xi^{i}_{\alpha}=$ $0\,.$ (2.11e)
@@
$\xi_{i}^{\alpha}=$ $\frac{1}{2}{\cal D}_{i}^{\alpha}\rho\,$, $\xi_{\alpha\dot{\beta}}=$ $\frac{i}{2S^{2}}S_{ij}{\cal D}^{i}_{\alpha}\bar{\cal D}^{j}_{\dot{\beta}}\rho\,$, (2.12)
$\lambda^{ij}=$ $\rho S^{ij}\,$, $\lambda_{\alpha\beta}=$ $\frac{1}{4}{\cal D}^{k}_{\alpha}{\cal D}_{\beta k}\rho\,.$ (2.13)
@@
$\left({\cal D}^{\alpha i}{\cal D}^{j}_{\alpha}+4S^{ij}\right)\rho=$ $0\,$, (2.14a)
$\left({\cal D}^{i}_{\alpha}\bar{\cal D}^{j}_{\dot{\beta}}-\frac{1}{2S^{2}}S^{ij}S_{kl}{\cal D}^{k}_{\alpha}\bar{\cal D}^{l}_{\dot{\beta}}\right)\rho=$ $0\,$, (2.14b)
${\cal D}_{a}\rho=$ $0\,.$ (2.14c)
@@
$\delta_{\xi}\Phi=\left(\xi+\frac{1}{2}\lambda^{ab}M_{ab}+\rho S^{ij}J_{ij}\right)\Phi\,.$ (2.15)
@@ 3 Introducing harmonics
$D^{++},D^{--}]=D^{0}\,,\quad[D^{0},D^{++}]=2D^{++}\,,\quad[D^{0},D^{--}]=-2D^{--}\,.$ (3.1)
With every iso-tensor superfield $\Phi^{i_{1}\ldots i_{q}}=\Phi^{(i_{1}\ldots i_{q})}$ is associated a harmonic superfield
$\Phi^{(q)}(z,u)=u^{+}_{i_{1}}\ldots u^{+}_{i_{q}}\Phi^{i_{1}\ldots i_{q}}(z)\,.$ (3.2)
By construction, this superfield obeys
$D^{++}\Phi^{(q)}=0\,,\qquad D^{0}\Phi^{(q)}=q\Phi^{(q)}\,.$ (3.3)
For these superfields, the $U(1)$ generator ${\cal J}:=S^{ij}J_{ij}$ works as follows:
$S^{ij}J_{ij}\Phi^{(q)}=(S^{--}D^{++}-S^{++}D^{--}+S^{+-}D^{0})\Phi^{(q)}\,$, (3.4)
where $S^{\pm\pm}=u^{\pm}_{i}u^{\pm}_{j}S^{ij}$. Eq. (2.15) implies the transformation law of the harmonic superfield $\Phi^{(q)}$ with no Lorentz indices under the isometry of AdS4|8:
$\delta_{\xi}\Phi^{(q)}=(\xi^{a}{\cal D}_{a}-\xi^{+\alpha}{\cal D^{-}_{\alpha}}+\xi^{-\alpha}{\cal D}^{+}_{\alpha}-\bar{\xi}^{-}_{\dot{\alpha}}\bar{\cal D}^{+\dot{\alpha}}+\bar{\xi}^{+}_{\dot{\alpha}}\bar{\cal D}^{-\dot{\alpha}}+\rho{\cal J})\Phi^{(q)}\,$, (3.5)
@@
For analytic superfields obeying ${\cal D}^{+}_{\alpha}\Phi_{A}^{(q)}=\bar{\cal D}^{+}_{\dot{\alpha}}\Phi_{A}^{(q)}=0$ and $D^{++}\Phi^{(q)}_{A}=0$ the variation (3.5) simplifies to
$\delta_{\xi}\Phi_{A}^{(q)}=(\xi^{a}{\cal D}_{a}-\xi^{+\alpha}{\cal D^{-}_{\alpha}}+\bar{\xi}^{+}_{\dot{\alpha}}\bar{\cal D}^{-\dot{\alpha}}-\rho S^{++}D^{--}+q\rho S^{+-})\Phi_{A}^{(q)}\,.$ (3.6)
The algebra of covariant derivatives (2.7) may be re-written in terms of covariant derivatives ${\cal D}^{\pm}_{\alpha}$ and $\bar{\cal D}^{\pm}_{\dot{\alpha}}$. Among these (anti)commutation relations we will need the following ones:
$\{{\cal D}^{+}_{\alpha},\bar{\cal D}^{-}_{\dot{\alpha}}\}=$ $-2i\sigma^{a}_{\alpha\dot{\alpha}}{\cal D}_{a}\,$, (3.7a)
$\{{\cal D}^{+}_{\alpha},{\cal D}^{+}_{\beta}\}$= $4S^{++}M_{\alpha\beta}\,$, $\{\bar{\cal D}^{+}_{\dot{\alpha}},\bar{\cal D}^{+}_{\dot{\beta}}\}$= $-4S^{++}\bar{M}_{\dot{\alpha}\dot{\beta}}\,$, (3.7b)
$\{{\cal D}^{+}_{\alpha},{\cal D}^{-}_{\beta}\}=$ $4S^{+-}M_{\alpha\beta}-2\varepsilon_{\alpha\beta}{\cal J}\,$, $\{\bar{\cal D}^{+}_{\dot{\alpha}},\bar{\cal D}^{-}_{\dot{\beta}}\}=$ $-4S^{+-}\bar{M}_{\dot{\alpha}\dot{\beta}}+2\varepsilon_{\dot{\alpha}\dot{\beta}}{\cal J}\,$, (3.7c)
$[{\cal D}_{a},{\cal D}^{\pm}_{\alpha}]=$ $\frac{i}{2}(\sigma_{a})_{\alpha\dot{\alpha}}(S^{\pm+}\bar{\cal D}^{-\dot{\alpha}}-S^{\pm-}\bar{\cal D}^{+\dot{\alpha}})\,$, (3.7d)
$[{\cal D}_{a},\bar{\cal D}^{\pm}_{\dot{\alpha}}]=$ $\frac{i}{2}(\sigma_{a})_{\alpha\dot{\alpha}}(S^{\pm+}{\cal D}^{-\alpha}-S^{\pm-}{\cal D}^{+\alpha})\,.$ (3.7e)
The anticommutators (3.7b) imply the following identities
${\cal D}^{+}_{\alpha}[({\cal D}^{+})^{2}+4S^{++}]\Phi^{(q)}=$ $0\,$, (3.8)
$\bar{\cal D}^{+}_{\dot{\alpha}}[(\bar{\cal D}^{+})^{2}+4S^{++}]\Phi^{(q)}=$ $0\,.$ (3.9)
@@
$\hat{\Box}:=\frac{1}{32}[({\cal D}^{+})^{2}+4S^{++}][(\bar{\cal D}^{+})^{2}+4S^{++}]D^{--}D^{--}$ (3.11)
preserves the analyticity and maps any analytic superfield to another analytic superfield. Using the algebra of covariant derivatives (3.7) it is possible to derive the following representation for this operator when it acts on an analytic superfield with $U(1)$ charge $q$:
$\hat{\Box}\Phi_{A}^{(q)}=$ $\Big[{\cal D}^{a}{\cal D}_{a}-\frac{1}{4}S^{++}({\cal D}^{-})^{2}-\frac{1}{4}S^{++}(\bar{\cal D}^{-})^{2}+\frac{1}{2}S^{++}S^{++}D^{--}D^{--}$
$+{\cal J}({\cal J}+S^{++}D^{--}-3S^{+-})\Big]\Phi_{A}^{(q)}\,$, (3.12)
This operator is an AdS4|8 counterpart of the flat harmonic superspace d’Alembert operator introduced in Ref. [32].
4 Analytic action principle

### v2

@@ Abstract
We develop an analytic action principle in $\mathcal{N}=2$ AdS4 harmonic superspace and apply it to study the component structure of the free $q$-hypermultiplet model.
I Introduction and conclusions
Superspace formulations of supersymmetric field theories provide explicit off-shell realizations of the Poincaré superalgebra, whose closure does not require the equations of motion. Such formulations are particularly advantageous for models with extended supersymmetry, since hypermultiplet matter theories require an infinite number of auxiliary fields off shell, which naturally emerge within the harmonic [1, 2] and projective [3, 4] superspace approaches. Being complementary [5], these two formalisms are especially useful for constructing classical actions of higher-spin field theories with extended supersymmetry [6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. Generalizing ${\cal N}=2$ supersymmetric field theory to the 4D anti de-Sitter background (AdS4), however, represents a nontrivial problem, as the corresponding geometry in the harmonic superspace has not yet been fully developed. Although general supergravity theories were formulated in 4D $\mathcal{N}=2$ harmonic superspace (see Refs. [1, 16, 17, 18, 19, 20, 21]), the specific features of ${\cal N}=2$ AdS4 harmonic superspace remain insufficiently explored, and the construction of classical actions for gauge and matter theories in these superspaces has not yet been properly addressed. In contrast, the analytic action principle in the AdS5 harmonic superspace was developed in the works [22, 23].
@@ I Introduction and conclusions
Despite the progress achieved in formulating field theories in ${\cal N}=2$ AdS4 harmonic superspace in the present paper and in Ref. [24], the problem remains far from complete. The approach of Ref. [24] is based on manifestly analytic superfields and employs the so-called analytic basis in superspace. In the present work, we use the central basis, in which analyticity is not manifest but is instead imposed through superfield constraints. This latter formulation is, however, fully covariant and explicitly geometric, owing to the developments in Refs. [25, 26, 27, 28, 29, 30, 31, 32] where AdS4|8 superspace was obtained as a particular solution of $\mathcal{N}=2$ superfield supergravity. In general $\mathcal{N}=2$ supergravity, these two coordinate bases are related by the so-called bridge superfield, originally constructed in Ref. [20]. The explicit construction of this bridge superfield for the ${\cal N}=2$ AdS4 harmonic superspace background remains an important open problem, which will be addressed elsewhere.
The rest of the paper is organized as follows. In the next section, we present some background information about the geometry of the AdS4|8 superspace and its isometry generated by the Killing vectors. In Sec. III we extend this superspace with harmonic variables and introduce analytic superspace projectors. In Sec. IV we show that these projectors may be used for construction a superfield action for an analytic Lagrangian in the harmonic superspace. We develop also the analytic action principle in a fully covariant and coordinate independent form. Then we apply this action principle for calculating the component structure of the classical action in the free $q$-hypermultiplet model. In our work, we follow superspace conventions adopted in Refs. [25, 26, 27].
II Isometries of AdS4|8 superspace
@@
$[M_{ab},{\cal D}_{c}]=$ $2\eta_{c[a}{\cal D}_{b]}\,$, (3a)
$[M_{\alpha\beta},{\cal D}^{i}_{\gamma}]=$ $\varepsilon_{\gamma(\alpha}{\cal D}^{i}_{\beta)}\,$, (3b)
$[\bar{M}_{\dot{\alpha}\dot{\beta}},\bar{\cal D}^{i}_{\dot{\gamma}}]=$ $\varepsilon_{\dot{\gamma}(\dot{\alpha}}\bar{\cal D}^{i}_{\dot{\beta})}\,.$ (3c)
@@
$M_{ab}=$ $(\sigma_{ab})^{\alpha\beta}M_{\alpha\beta}-(\tilde{\sigma}_{ab})^{\dot{\alpha}\dot{\beta}}\bar{M}_{\dot{\alpha}\dot{\beta}}\,$, (4a)
$M_{\alpha\beta}=$ $\frac{1}{2}(\sigma^{ab})_{\alpha\beta}M_{ab}\,$, (4b)
$\bar{M}_{\dot{\alpha}\dot{\beta}}=$ $-\frac{1}{2}(\tilde{\sigma}^{ab})_{\dot{\alpha}\dot{\beta}}M_{ab}\,.$ (4c)
@@
$J_{kl},{\cal D}^{i}_{\alpha}]=-\delta^{i}_{(k}{\cal D}_{\alpha l)}\,,\qquad[J_{kl},\bar{\cal D}_{i}^{\dot{\alpha}}]=-\varepsilon_{i(k}\bar{\cal D}^{\dot{\alpha}}_{l)}\,.$ (6)
@@
$\{{\cal D}^{i}_{\alpha},{\cal D}^{j}_{\beta}\}=$ $4S^{ij}M_{\alpha\beta}+2\varepsilon_{\alpha\beta}\varepsilon^{ij}S^{kl}J_{kl}\,$, (7a)
$\{\bar{\cal D}_{\dot{\alpha}i},\bar{\cal D}_{\dot{\beta}j}\}=$ $-4S_{ij}\bar{M}_{\dot{\alpha}\dot{\beta}}+2\varepsilon_{\dot{\alpha}\dot{\beta}}\varepsilon_{ij}S^{kl}J_{kl}\,$, (7b)
$\{{\cal D}^{i}_{\alpha},\bar{\cal D}_{j}^{\dot{\beta}}\}=$ $-2i\delta^{i}_{j}(\sigma^{c})_{\alpha}{}^{\dot{\beta}}{\cal D}_{c}\,$, (7c)
$[{\cal D}_{a},{\cal D}^{j}_{\beta}]=$ $\frac{i}{2}(\sigma_{a})_{\beta\dot{\gamma}}S^{jk}\bar{\cal D}^{\dot{\gamma}}_{k}\,$, (7d)
$[{\cal D}_{a},\bar{\cal D}_{j}^{\dot{\beta}}]=$ $\frac{i}{2}(\sigma_{a})^{\alpha\dot{\beta}}S_{jk}{\cal D}^{k}_{\alpha}\,$, (7e)
$[{\cal D}_{a},{\cal D}_{b}]=$ $-S^{2}M_{ab}\,$, (7f)
@@
${\cal D}_{A}S^{ij}=0\,,\qquad\overline{S^{ij}}=S_{ij}\,.$ (8)
In Ref. [25], it was shown that one can impose a special gauge on the $SU(2)$ connection such that the torsion $S^{ij}$ becomes a constant, $S^{ij}=c^{ij}=\mathrm{const}$. These constants $c^{ij}$ play a central role in the construction of the AdS4 harmonic superspace in Ref. [24]. In the present work, we do not impose this gauge and therefore keep the torsion $S^{ij}$ covariantly constant.
@@
$\xi=\xi^{A}{\cal D}_{A}=\xi^{a}{\cal D}_{a}+\xi^{\alpha}_{i}{\cal D}^{i}_{\alpha}+\bar{\xi}^{i}_{\dot{\alpha}}\bar{\cal D}_{i}^{\dot{\alpha}}$ (9)
@@
$\left[\delta_{\xi},{\cal D}_{A}\right]=\left[\xi+\frac{1}{2}\lambda^{ab}M_{ab}+\lambda^{ij}J_{ij},{\cal D}_{A}\right]=0\,$, (10)
for some real antisymmetric tensor $\lambda^{cd}(z)=-\lambda^{dc}(z)$ and real symmetric tensor $\lambda^{ij}=\overline{\lambda_{ij}}$.
In Ref. [25], it was shown that the parameters $\lambda^{ij}$ may be expressed via a single superfield $\rho$. Indeed, $S^{ij}$ being a component of supertorsion should be invariant under the isometry transformation, $\left[\delta_{\xi},S^{ij}\right]=0$. Since in the AdS4|8 superspace this supertorsion is covariantly constant, see Eq. (8), we find that $S^{ij}$ is invariant under the $SU(2)$ action of $\lambda^{ij}J_{ij}$, which implies $\lambda^{ij}=\rho S^{ij}$. This superfield, as well as other components of the Killing vector $\xi^{A}$ obey a number of constraints which follow from Eq. (10):
${\cal D}^{i}_{\alpha}\xi^{\beta}_{j}-\rho S^{i}{}_{j}\delta_{\alpha}^{\beta}-\lambda_{\alpha}{}^{\beta}\delta^{i}_{j}=$ $0\,$, (11a)
$\bar{\cal D}_{i}^{\dot{\alpha}}\xi^{\beta}_{j}-\frac{i}{2}S_{ij}\xi^{\beta\dot{\alpha}}=$ $0\,$, (11b)
$\bar{\cal D}_{i}^{\dot{\alpha}}\xi^{b}+2i(\sigma^{b})_{\beta}{}^{\dot{\alpha}}\xi_{i}^{\beta}=$ $0\,$, (11c)
${\cal D}^{i}_{\alpha}\lambda^{cd}-4S^{ij}\xi_{j}^{\beta}(\sigma^{cd})_{\alpha\beta}=$ $0\,$, (11d)
${\cal D}^{i}_{\alpha}\rho-2\xi^{i}_{\alpha}=$ $0\,.$ (11e)
@@
$\xi_{i}^{\alpha}=$ $\frac{1}{2}{\cal D}_{i}^{\alpha}\rho\,$, $\xi_{\alpha\dot{\beta}}=$ $\frac{i}{2S^{2}}S_{ij}{\cal D}^{i}_{\alpha}\bar{\cal D}^{j}_{\dot{\beta}}\rho\,$, (12)
$\lambda^{ij}=$ $\rho S^{ij}\,$, $\lambda_{\alpha\beta}=$ $\frac{1}{4}{\cal D}^{k}_{\alpha}{\cal D}_{\beta k}\rho\,.$ (13)
@@
$\left({\cal D}^{\alpha i}{\cal D}^{j}_{\alpha}+4S^{ij}\right)\rho=$ $0\,$, (14a)
$\left({\cal D}^{i}_{\alpha}\bar{\cal D}^{j}_{\dot{\beta}}-\frac{1}{2S^{2}}S^{ij}S_{kl}{\cal D}^{k}_{\alpha}\bar{\cal D}^{l}_{\dot{\beta}}\right)\rho=$ $0\,$, (14b)
${\cal D}_{a}\rho=$ $0\,.$ (14c)
@@
$\delta_{\xi}\Phi=\left(\xi+\frac{1}{2}\lambda^{ab}M_{ab}+\rho S^{ij}J_{ij}\right)\Phi\,.$ (15)
@@ III Introducing harmonics
$[D^{++},D^{--}]=$ $D^{0}\,$, (16a)
$[D^{0},D^{++}]=$ $2D^{++}\,$, (16b)
$[D^{0},D^{--}]=$ $-2D^{--}\,.$ (16c)
Following the routine of the harmonic superspace approach [1], with every iso-tensor superfield with no Lorentz indices
$\Phi^{i_{1}\ldots i_{m+n}}(z)=\Phi^{(i_{1}\ldots i_{m+n})}(z)\,$, (17)
we associate a harmonic superfield
$\Phi^{(m,n)}(z,u)=u^{+}_{i_{1}}\ldots u^{+}_{i_{m}}u^{-}_{i_{m+1}}\ldots u^{-}_{i_{m+n}}\Phi^{i_{1}\ldots i_{m+n}}(z)\,.$ (18)
By construction this superfield has $U(1)$ charge q=mnq=m-n,
$D^{0}\Phi^{(m,n)}=(m-n)\Phi^{(m,n)}\,$, (19)
and, thus, it is a smooth function over the homogeneous space $S^{2}\simeq SU(2)/U(1)$. More generally, given a set of superfields with arbitrary number of $SU(2)$ indices of the type (17), we define a general harmonic superfield with a given $U(1)$ charge $q$:
[…]
$D^{0}\Phi^{(q)}=q\Phi^{(q)}\,.$ (21)
By construction, the harmonic variables are invariant under the AdS4|8 isometry,
$\delta_{\xi}u^{\pm}_{i}=0\,.$ (22)
This means that they are inert under the $SU(2)$ generator $J_{ij}$. As a consequence, the $U(1)$ generator $S^{ij}J_{ij}$ acting on the harmonic superfields (20) is represented by the following operator:
${\cal J}\Phi^{(q)}:=$ $S^{ij}J_{ij}\Phi^{(q)}$
$=$ $(S^{--}D^{++}-S^{++}D^{--}+S^{+-}D^{0})\Phi^{(q)}\,$, (23)
where
$S^{\pm\pm}=u^{\pm}_{i}u^{\pm}_{j}S^{ij}\,.$ (24)
Eq. (15) implies the transformation law of the harmonic superfield $\Phi^{(q)}$ with no Lorentz indices under the isometry of AdS4|8:
$\delta_{\xi}\Phi^{(q)}=$ $(\xi^{a}{\cal D}_{a}-\xi^{+\alpha}{\cal D^{-}_{\alpha}}+\xi^{-\alpha}{\cal D}^{+}_{\alpha}$
$-\bar{\xi}^{-}_{\dot{\alpha}}\bar{\cal D}^{+\dot{\alpha}}+\bar{\xi}^{+}_{\dot{\alpha}}\bar{\cal D}^{-\dot{\alpha}}+\rho{\cal J})\Phi^{(q)}\,$, (25)
@@
For analytic superfields obeying
${\cal D}^{+}_{\alpha}\Phi_{A}^{(q)}=\bar{\cal D}^{+}_{\dot{\alpha}}\Phi_{A}^{(q)}=0\,$, (26)
the variation (25) simplifies to
$\delta_{\xi}\Phi_{A}^{(q)}=(\xi^{a}{\cal D}_{a}-\xi^{+\alpha}{\cal D^{-}_{\alpha}}+\bar{\xi}^{+}_{\dot{\alpha}}\bar{\cal D}^{-\dot{\alpha}}+\rho{\cal J})\Phi_{A}^{(q)}\,.$ (27)
Note that the analyticity constraint (26) is preserved by the isometry transformations in AdS4|8, ${\cal D}^{+}_{\alpha}\delta_{\xi}\Phi_{A}^{(q)}=\bar{\cal D}^{+}_{\dot{\alpha}}\delta_{\xi}\Phi_{A}^{(q)}=0$. This property follows from the definition of the Killing vector (10) and the condition (22).
The algebra of covariant derivatives (7) may be re-written in terms of covariant derivatives ${\cal D}^{\pm}_{\alpha}$ and $\bar{\cal D}^{\pm}_{\dot{\alpha}}$. Among these (anti)commutation relations we will need the following ones:
$\{{\cal D}^{+}_{\alpha},\bar{\cal D}^{-}_{\dot{\alpha}}\}=$ $-2i\sigma^{a}_{\alpha\dot{\alpha}}{\cal D}_{a}\,$, (28a)
$\{{\cal D}^{+}_{\alpha},{\cal D}^{+}_{\beta}\}$= $4S^{++}M_{\alpha\beta}\,$, (28b)
$\{\bar{\cal D}^{+}_{\dot{\alpha}},\bar{\cal D}^{+}_{\dot{\beta}}\}$= $-4S^{++}\bar{M}_{\dot{\alpha}\dot{\beta}}\,$, (28c)
$\{{\cal D}^{+}_{\alpha},{\cal D}^{-}_{\beta}\}=$ $4S^{+-}M_{\alpha\beta}-2\varepsilon_{\alpha\beta}{\cal J}\,$, (28d)
$\{\bar{\cal D}^{+}_{\dot{\alpha}},\bar{\cal D}^{-}_{\dot{\beta}}\}=$ $-4S^{+-}\bar{M}_{\dot{\alpha}\dot{\beta}}+2\varepsilon_{\dot{\alpha}\dot{\beta}}{\cal J}\,$, (28e)
$[{\cal D}_{a},{\cal D}^{\pm}_{\alpha}]=$ $\frac{i}{2}(\sigma_{a})_{\alpha\dot{\alpha}}(S^{\pm+}\bar{\cal D}^{-\dot{\alpha}}-S^{\pm-}\bar{\cal D}^{+\dot{\alpha}})\,$, (28f)
$[{\cal D}_{a},\bar{\cal D}^{\pm}_{\dot{\alpha}}]=$ $\frac{i}{2}(\sigma_{a})_{\alpha\dot{\alpha}}(S^{\pm+}{\cal D}^{-\alpha}-S^{\pm-}{\cal D}^{+\alpha})\,.$ (28g)
The anticommutators (28b,28c) imply the following identities for a general unconstrained harmonic superfield of the form (20)
${\cal D}^{+}_{\alpha}[({\cal D}^{+})^{2}+4S^{++}]\Phi^{(q)}=$ $0\,$, (29)
$\bar{\cal D}^{+}_{\dot{\alpha}}[(\bar{\cal D}^{+})^{2}+4S^{++}]\Phi^{(q)}=$ $0\,.$ (30)
@@
$\hat{\Box}:=\frac{1}{32}[({\cal D}^{+})^{2}+4S^{++}][(\bar{\cal D}^{+})^{2}+4S^{++}]D^{--}D^{--}$ (32)
preserves the analyticity and maps any analytic superfield to another analytic superfield. Using the algebra of covariant derivatives (28) it is possible to derive the following representation for this operator when it acts on an analytic superfield with $U(1)$ charge $q$:
$\hat{\Box}\Phi_{A}^{(q)}=$ $\Big[{\cal D}^{a}{\cal D}_{a}-\frac{1}{4}S^{++}({\cal D}^{-})^{2}-\frac{1}{4}S^{++}(\bar{\cal D}^{-})^{2}$
$+\frac{1}{2}S^{++}S^{++}D^{--}D^{--}$
$+{\cal J}({\cal J}+S^{++}D^{--}-3S^{+-})\Big]\Phi_{A}^{(q)}\,.$ (33)
This operator is an AdS4 counterpart of the flat ${\cal N}=2$ harmonic superspace d’Alembert operator introduced in Ref. [33].
IV Analytic action principle